OpenStudy (hitaro9):

A water trough is 13 feet long, and its cross section is an equilateral triangle with sides 4 feet long. Water is pumped into the trough at a rate of 10 cubic feet per second. What is the height h of an equilateral triangle of side length s? How fast is the water level rising when the depth of the water is 1/2 foot?

OpenStudy (hitaro9):

I'm kinda lost as to what this problem is even supposed to look like.

zepdrix (zepdrix):

I think this might be the shape we're looking for.

zepdrix (zepdrix):

They give us some information: \(\Large V'=10\), The rate at which water is coming in. And they want us to find \(\Large h'\), the rate at which the water is rising, when \(\Large h=\dfrac{1}{2}\).

zepdrix (zepdrix):

Hmm...

OpenStudy (wolf1728):

Height of the trough² is 4² - 2² = 12 trough height = sqrt(12) = 2*sqrt(3) = 3.464